# Krylov Complexity in early universe

### v2

@@ 3.2 Various inflationary potentials
where $M_{p}=(8\pi G)^{-\frac{1}{2}}$ is the definition of the reduced Planck mass. Additionally, $M$ represents the energy scale, and its value can be determined based on the amplitude of the observed power spectrum for primordial perturbations. Being armed with the solution of Eq. (20) and $dt=ad\eta$, we could rewrite $V_{\phi\phi}$ as
@@
$\begin{cases}&\ V_{1,\phi\phi}=m^{2},\\ &\ V_{2,\phi\phi}=M^{2}\Bigl{(}-e^{-\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}+2e^{-2\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}\Bigr{)},\\ &\ V_{3,\phi\phi}=m^{2}+3\lambda A^{2}a^{-3}\sin^{2}(m_{\rm eff}a\eta)\end{cases}$ (23)
where $\phi(t)=Aa^{-\frac{3}{2}}\sin(m_{\text{eff}}t)$. For the effective mass $m_{\text{eff}}$, we have $m_{\text{eff}}^{2}=\frac{d^{2}V(x)}{dx^{2}}$, and it is evident that $m_{\text{eff}}=\sqrt{V_{i,\phi\phi}}$ in our work $(i=1,2,3)$. According to effective mass for the distinctive inflationary models, we need the iterations for $V_{2}$ and $V_{3}$ in terms of $\log_{10}a$ followed by Li:2024kfm . To determine whether the functions $V_{2}$ and $V_{3}$ converge within the defined interval, we performed iterations at multiple sampling points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ with the step $0.05$. Our findings reveal that all points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ are all convergent.
@@
Figure 1: The numeric of $V_{,\phi\phi}$ at first 50 iterations of potential $V_{2}$ and $V_{3}$ in $\log_{10}a=0.2$ and $\log_{10}a=2.8$, in which we set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
@@
As an illustration, Fig. 1 clearly shows the iterations for potential $V_{2}$ and $V_{3}$, where we have performed $50$ times iterations for two specific values of $\log_{10}a$. Let us mention the range of parameters as shown in Fig. 1, the potential can be constraint by the COBE normalization Planck:2015sxf , whose values are bounded below $10^{-10}$ in Planck units, and the inflationary field is a large field and its maximal value is around $100$. Consequently, the effective mass for various potentials is around $10^{-6}$ and other parameters are set according to this logic.
@@
Figure 2: The numeric of $V_{,\phi\phi}$ in terms of $\log_{10}a$ for potential $V_{2}$, and $V_{3}$ in RD and MD, in which we replace its exact value with the third iteration of $V_{2}$ and $V_{3}$. We set $\theta_{0}=0$, $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
@@
Fig. 2 depicts the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ in RD and MD, using the third iteration of $V_{2}$ and $V_{3}$ via Fig. 1.The evolution of inflation is neglected due to the violation of slow-roll conditions. From Fig. 2, we could see that the effective mass of $R^{2}$ inflation will decrease in RD and then it will grow to the constant ($10^{-12}$) with the oscillations. For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants. But, it indicates that the effective mass of $V_{2}$ and $V_{3}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
@@
In this section, we provide an introduction to inflation, RD, and MD, which are characterized by the corresponding scale factor. We then present three essential inflationary potentials that encompass most types of inflationary models. Additionally, we provide an overview of preheating, during which the background of the inflaton field oscillates around the minimum. Subsequently, we incorporate the background solution for the inflaton into the numerical calculation of the effective mass for these three distinct potentials.
4 Evolution of two-mode squeezed state
In this section, we will follow the method of Li:2024kfm for investigating the evolution of $r_{k}$ and $\phi_{k}$ in inflation, RD and MD. We have evaluated the quantum circuit complexity in various inflationary frameworks Liu:2021nzx ; Li:2021kfq ; Li:2023ekd .
4.1 Two-mode squeezed state
@@
For obtaining the two-mode squeezed state, we should introduce the unitary operator as follows,
@@
^$\hat{\mathcal{R}}_{k}(\theta_{k})=\exp\bigg{[}-i\theta_{k}(\eta)\big{(}\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}\big{)}\bigg{]}$ (26)
where $\hat{\mathcal{S}}_{k}$ is the squeezed parameter and ^$\hat{\mathcal{R}}_{k}(\theta_{k})$ is the rotation parameter that will be neglected since it only contributes the phase factor of wave function. Then, by applying $\hat{S}_{\vec{k}}(r_{k},\phi_{k})$ to the vacuum state, we can express the wave function of the two-mode squeezed state as follows
@@
Being armed with this wave function, we could proceed the numeric of $\phi_{k}$ and $r_{k}$. But the evolution of $\phi_{k}$ and $r_{k}$ is based on the total action (19).
4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
In this work, we will evaluate the impact of potential for the Krylov complexity, where the Mukhanov-Sasaki variable contains the contribution of potential. Consequently, we cannot observe the explicit influence from potential. Therefore, we will directly perturb the total action into the second order,
[…]
@@ 4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
According to Eq. (28) and Eq. (29), we could obtain
[…]
@@
we could obtain the evolution of equation of $r_{k}$ and $\phi_{k}$
@@
$\begin{split}&r_{k}^{{}^{\prime}}=-\frac{a^{\prime}}{a}\cos(2\phi_{k}),\\ &\phi_{k}^{{}^{\prime}}=-k+\frac{a^{\prime}}{a}\sin(2\phi_{k})\coth(2r_{k}).\end{split}$ (36)
It is very difficult to obtain the numeric of Eq. (35), thus we could make a variable transformation, $y=\log_{10}a$, to numerically solve Eq. (35),
@@
Figure 3: The numerical solution of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD, $k=0.005$ at MD for simplicity. The effect of potential is always ignored in slow-roll inflation. But in RD and MD periods, the impact of potential is taken into account.
Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$. Therefore, we only provide the numerical values of $r_{k}$ in Fig. 3 based on Eq. (37). In this equation, we have set $H_{0}=1$ for simplicity and used various values of $k$ in different time periods. The first panel of Fig. 3 illustrates the evolution of $r_{k}$ during inflation. There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm . As we transit to RD, we observe that $r_{k}$ will increase to some constant value that oscillates around this value. Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$. This observation also applies to MD, where the evolution trend is similar to RD. However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases. Although we cannot distinguish between the various potentials using Fig. 3, our numerical data still demonstrates that the evolution of $r_{k}$ is significantly different whether there is a potential or not. It’s worth noting that we cannot provide the values of $r_{k}$ without a potential, as this is not physically meaningful. The potential cannot be ignored due to the violation of slow-roll conditions.
In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within a single-field inflationary framework. Our research indicates that the value of $r_{k}$ increases during inflation. The behavior of $r_{k}$ in RD and MD eras is similar and will oscillate around constant values. Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$. Additionally, we have observed that the scale factor and momentum have a significant impact on the evolution of $r_{k}$. All of the details are encoded in Fig. 3.
5 Krylov complexity in closed system
In this section, we will investigate the Krylov complexity in the method of closed system Liu:2021nzx ; Li:2021kfq ; Li:2023ekd , consisting of the Krylov complexity and Krylov entropy.
5.1 Krylov complexity
@@
For obtaining the Krylov complexity, the most essential quantity is the wave function. First, the two-mode squeezed state wave function can be represented by Krylov basis as follows,
$\mathcal{O}(t)=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (38)
Then, we could find the explicit wave function $\phi_{n}$
@@
We can define the Krylov compplexity in terms of $\phi_{n}$
@@
It is explicitly showing that the Krylov complexity is independent of $\phi_{k}$, that is why we only list the figure of $r_{k}$.
@@
Being armed with Eq. (40), we could give its numeric in Fig. 4, indicating the evolution of Krylov complexity in inflation, RD, and MD. Fig. 4 clearly shows that the Krylov complexity is enhanced in inflation which is consistent with the standard case of the modified dispersion within the framework Li:2024kfm . Similar work investigated the Krylov complexity for the exponential expansion for de-Sitter spacetime corresponding to the inflation Adhikari:2022oxr . Ref. Adhikari:2022oxr mentioned that the Krylov complexity equals the average particle number proportional to the volume. Thus we could see that this simple analysis is consistent with the CV conjuncture Stanford:2014jda , in which the complexity is equaled to the volume of wormhole under the framework of holography. However, the method of Adhikari:2022oxr is for the closed system.
In the case of RD and MD, it is observed that the Krylov complexity will increase to certain constant values that fluctuate around this value. Additionally, different patterns of inflationary models tend to produce similar trends of Krylov complexity in different periods. Our analysis aligns with the findings of Barbon:2019wsy , where the Krylov complexity will saturate some constant values even after the scrambling time in a closed system. The scrambling time will be assessed in the context of an open system, particularly with regard to the Lanczos coefficient, Lyapunov index, $\it e.t.c$.
5.2 Krylov Entropy
In a quantum system, entropy is defined as a measure of the disorder of a system. The Krylov entropy (K-entropy) can be used to test the level of disorder of curvature perturbation. Following Barbon:2019wsy , the K-entropy has the following form

### v3

@@ III.2 Various inflationary potentials
where $M_{p}=(8\pi G)^{-\frac{1}{2}}$ is the reduced Planck mass. Here, $M$ denotes an energy scale whose value is fixed by the observed amplitude of the primordial power spectrum. Using the solution of Eq. (20) and the relation $dt=ad\eta$, we can express $V,_{\phi\phi}$ as
@@
$\begin{cases}&\ V_{\rm chaotic,\phi\phi}=m^{2},\\ &\ V_{\rm R^{2},\phi\phi}=M^{2}\Bigl(-e^{-\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}+2e^{-2\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}\Bigr),\\ &\ V_{\rm Higss,\phi\phi}=m^{2}+3\lambda A^{2}a^{-3}\sin^{2}(m_{\rm eff}a\eta)\end{cases}$ (24)
where $\phi(t)=Aa^{-\frac{3}{2}}\sin(m_{\text{eff}}t)$ and we have used $dt=ad\eta$. The definition of effective mass in Eq. (21) has been employed, and clearly $m_{\text{eff}}=\sqrt{V_{i,\phi\phi}}$ in our work ($\rm i=chaotic,R^{2},Higgs$). To simplify the subsequent numerical simulations, we define the variable $y=\log_{10}a$. To numerically obtain the effective mass of the distinct inflationary models $V_{\rm R^{2}}$ and $V_{\rm Higgs}$, iterative evaluations in terms of $\log_{10}a$ are required, based on Eq. (24). Let us clarify the concept of iterations for the effective mass. The Newton-Raphson method is adopted for its computational efficiency and fast convergence rate. In (24), the effective mass $m_{\rm eff}$ defined by (21) carries the same physical meaning as the left-hand side of (24). Thus, the relation (24) naturally lends itself to an iterative scheme, requiring an initial value $m_{0}$ for the effective mass. This value ($m_{0}^{2}=10^{-12}$) can be fixed by the COBE normalization Planck:2015sxf . In our implementation, we perform 50 iterations for the effective mass, meaning that for each single value of $y$, the calculation is repeated 50 times.
To check the convergence of $V_{\rm R^{2},\phi\phi}$ and $V_{\rm Higgs,\phi\phi}$ within the defined interval, we performed iterations at multiple sampling points from $\log_{10}a=0$ to $\log_{10}a=6$ with a step size of $0.05$ (this range covers from RD to MD). We find that all points in the range $\log_{10}a=0$ to $\log_{10}a=6$ are convergent.
@@
Figure 1: The numeric of $V_{,\phi\phi}$ at first 50 iterations of $V_{\rm R^{2},\phi\phi}$ and $V_{\rm Higgs,\phi\phi}$ in $\log_{10}a=0.2$ and $\log_{10}a=2.8$, in which we set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{\rm R^{2}}$, and $\lambda A^{2}=10^{-4}$ at $V_{\rm Higss}$.
@@
As an illustration, Fig. 1 shows the iterative results for $V_{\rm R^{2},\phi\phi}$ and $V_{\rm Higgs,\phi\phi}$, where 50 iterations were performed for two specific values of $\log_{10}a$. The parameters used in the figure are chosen as follows: the potential amplitude is constrained by the COBE normalization Planck:2015sxf , which bounds its value below $10^{-10}$ in Planck units, while the inflationary field is a large field with a maximum value around $100$ in these units. Consequently, the effective mass for the various potentials is of order $10^{-6}$, and all other parameters are set consistently with this scaling.
Let us now discuss the effects of the distinctive parameters of the various potentials. For $V_{\text{chaotic}}$, its effective mass is explicitly set by the inflaton mass, whose value can be fixed via the COBE normalizations. As for $V_{R^{2}}$, it contains one parameter $M$, which acts as the amplitude of the effective mass in Eq. (24); consequently, the range of $M$ is also constrained by the COBE normalization. Finally, the Higgs potential $V_{\text{Higgs}}$ involves three parameters: the Higgs mass $m$, the coupling constant $\lambda$, and the amplitude of the inflaton during preheating $A$. Again, using the COBE normalizations, we can determine the upper bounds on $m^{2}$ and $\lambda A^{2}$; values exceeding these orders would violate the observational constraints. Thus, appropriate parameter values can be set from observational data. As illustrated in Fig. 2, the effective masses for all the considered potentials lie within $10^{-12}$ in Planck units.
@@
Figure 2: The numeric of $V_{,\phi\phi}$ in terms of $\log_{10}a$ for potential $V_{\rm R^{2}}$, and $V_{\rm Higgs}$ in RD and MD, in which we replace its exact value with the iteration of $V_{\rm R^{2}}$, and $V_{\rm Higgs}$. We set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ for $V_{\rm R^{2}}$, and $\lambda A^{2}=10^{-4}$ for $V_{\rm Higgs}$.
@@
Fig. 2 shows the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ during the RD and MD eras, based on the iterative results for $V_{\rm R^{2}}$ and $V_{\rm Higgs}$ from Fig. 1. The evolution of the inflationary phase itself is neglected here due to the violation of slow-roll conditions in these periods. From Fig. 2, we observe that the effective mass for $R^{2}$ inflation ($V_{2}$) decreases during RD and then grows, with oscillations, toward a constant value of approximately $10^{-12}$. The behavior for the Higgs potential ($V_{\rm Higgs}$) is different: it exhibits a peak during RD, after which it decreases and settles into an approximately constant value. Notably, the effective masses for both $V_{\rm R^{2}}$ and $V_{\rm Higgs}$ are of nearly the same order of magnitude, consistent with the COBE normalization constraint Planck:2015sxf .
@@
In this section, we introduce the inflationary, RD, and MD eras, each described by its corresponding scale factor. We also present three representative inflationary potentials that cover a broad range of model types. An overview of preheating is given, during which the inflaton background oscillates around its potential minimum. Finally, we incorporate this background solution into a numerical evaluation of the effective mass for each of the three potentials.
IV Evolution of two-mode squeezed state
As we know, the curvature perturbation will become classical as the horizon exits. Ref. Grishchuk:1990bj already shown the amplification from quantum level to classical level during inflation is naturally a quantum squeezing process. Thus, the two-mode squeezed state is quite a natural choice for investigating the early universe. In this section, we will follow the method of Li:2024kfm to investigate the evolution of $r_{k}$ and $\phi_{k}$ during inflation, RD and MD.
IV.1 Two-mode squeezed state
@@
First, we will simply recall the two-mode squeezed state. To obtain the two-mode squeezed state, we introduce the unitary operator as follows,
@@
$\hat{\mathcal{R}}_{k}(\theta_{k})=\exp\bigg[-i\theta_{k}(\eta)\big(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}\big)\bigg]$ (27)
rwhere $\hat{\mathcal{S}}_{k}$ is the squeeze operator and $\hat{\mathcal{R}}_{k}(\theta_{k})$ is the rotation operator, which can be neglected as it only contributes an overall phase factor to the wave function. Applying $\hat{S}_{{k}}(r_{k},\phi_{k})$ to the vacuum state then yields the wave function of the two-mode squeezed state:
@@
With this wave function, we can now compute the evolution of the squeezed state, which is governed by the parameters $r_{k}$ and $\phi_{k}$. The derivation of the two-mode squeezed state in Eq. (28) is model-independent, involving only the action of the squeeze operator on the vacuum. Model-specific information is encoded in the parameters $r_{k}$ and $\phi_{k}$, whose evolution is determined by the Hamiltonian. This will be confirmed in later sections. Furthermore, the two-mode quantum state $|n;n\rangle_{\vec{k},-\vec{k}}$ is a Fock state, and its form is given by
$|n;n\rangle_{\vec{k},-\vec{k}}=\bigg[\frac{1}{n!}(c_{\vec{k}}^{\dagger}c_{-\vec{k}}^{\dagger})|0,0\rangle_{\vec{k},-\vec{k}}\bigg],$ (29)
thus, it automatically satisfies both orthogonality and normalization conditions. Here, we need to specify the initial Krylov operator $\mathcal{O}_{0}$ in our setup. As mentioned, the wave function is a two-mode squeezed state (28); the initial operator must satisfy the Lanczos algorithm given in Eq. (5). Following Ref. Adhikari:2022oxr , we choose $|\mathcal{O}_{0})=|0,0\rangle_{\vec{k},-\vec{k}}$ as our initial operator. This choice naturally fulfills the requirements of the Lanczos algorithm, which will be verified in Sec. VI.
IV.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
Numerical simulations of $\phi_{k}$ and $r_{k}$ require their evolution equations, which are derived from the Hamiltonian within the two-mode squeezed state formalism. This work focuses on evaluating the impact of the potential on Krylov complexity. The conventional Mukhanov-Sasaki variable is not ideal for this purpose, as it encapsulates the combined effect of the potential and other factors. To explicitly track the potential’s role, we instead perform a direct second-order perturbation of the total action.
$S=\frac{1}{2}\int d\eta d^{3}x\Bigl[f^{\prime 2}-(\partial_{i}f)^{2}+(\frac{a^{\prime}}{a})^{2}f^{2}-2f^{\prime}f\frac{a^{\prime}}{a}-a^{2}V_{,\phi\phi}f^{2}\Bigr]$ (30)
where $f(\eta,\vec{x})$ is the inflaton perturbation defined via $\phi(\eta,\vec{x})=\bar{\phi}(\eta)+\frac{f(\eta,\vec{x})}{a(\eta)}$, with $\frac{f(\eta,\vec{x})}{a(\eta)}\ll 1$ in Planck units and this perturbation theory will be violated in strong gravitational system. In the current investigations, we focus on the perturbation of inflaton, which is responsible for the formation of large-scale structure; for further details, we refer the reader to Baumann:2009ds . Under the slow-roll condition during inflation, we neglect the potential term, i.e., we assume $\left(\frac{a^{\prime}}{a}\right)^{2}f^{2}-2f^{\prime}f\frac{a^{\prime}}{a}\gg a^{2}V_{,\phi\phi}f^{2}.$ The Hamiltonian is then constructed as $H=\int d^{3}x\,d\eta\,(\pi f^{\prime}-\mathcal{L})$, where $\mathcal{L}$ denotes the Lagrangian and $\pi$ is the conjugate momentum defined by
@@ IV.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
According to Eq. (30) and Eq. (31), we obtain
$H=\frac{1}{2}\int d^{3}xd\eta[\pi^{2}+(\partial_{i}f)^{2}+\frac{a^{\prime}}{a}(\pi f+f\pi)+a^{2}V_{,\phi\phi}f^{2}].$ (32)
@@
we obtain the evolution of equation of $r_{k}$ and $\phi_{k}$
@@
$\begin{split}&r_{k}^{{}^{\prime}}=-\frac{a^{\prime}}{a}\cos(2\phi_{k}),\\ &\phi_{k}^{{}^{\prime}}=-k+\frac{a^{\prime}}{a}\sin(2\phi_{k})\coth(2r_{k}),\end{split}$ (38)
where it is consistent with Bhattacharyya:2020rpy for $\frac{z^{\prime}}{z}=\frac{a^{\prime}}{a}$, since $\epsilon$ is constant. Directly solving Eq. (37) numerically is very difficult. Therefore, we introduce the variable transformation $y=\log_{10}a$ and solve Eq. (37) numerically in terms of $y$,
@@
Figure 3: The numerical solution of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ for RD, $k=0.005$ for MD. The effect of potential is always ignored in slow-roll inflation. But in RD and MD periods, the impact of potential is taken into account.
Before analyzing the figure for $r_{k}$, it is important to note that the Krylov complexity depends solely on this quantity. Therefore, in Fig. 3, we show only the numerical results for $r_{k}$ based on Eq. (39). For simplicity, we set $H_{0}=1$ and examine different values of the comoving wavenumber $k$ across various cosmological epochs.
The first panel of Fig. 3 illustrates the evolution of $r_{k}$ during inflation. Prior to horizon exit, $r_{k}$ exhibits oscillatory behavior, after which it grows linearly, consistent with the standard result reported in Li:2024kfm . Following the transition to the RD era, $r_{k}$ rises and then saturates, fluctuating around a nearly constant value. Notably, different inflationary potentials yield nearly identical trends for $r_{k}$, indicating that the specific form of the potential does not significantly influence its evolution.
A similar trend is observed during the MD era, where the behavior of $r_{k}$ closely resembles that in the RD era. However, we find that the value of the wavenumber $k$ has a pronounced effect: as $k$ decreases, $r_{k}$ displays increasingly prominent oscillations. We note that presenting results in RD and MD, the contribution of potential is taken into account since the violation of slow roll condtions.
In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within the single-field inflationary model. Our results indicate that $r_{k}$ grows during the inflationary epoch. In both the RD and MD eras, $r_{k}$ exhibits a similar qualitative behavior, saturating and then oscillating around an approximately constant value. Numerical analysis further shows that various inflationary models yield qualitatively similar trends for $r_{k}$. Additionally, we find that both the scale factor and the comoving wavenumber significantly influence the evolution of $r_{k}$. These features are summarized visually in Fig. 3.
V Krylov complexity in closed-system methodology
In this section, we conduct numerical simulations of Krylov complexity using the closed-system methodology Liu:2021nzx ; Li:2021kfq ; Li:2023ekd . This framework encompasses both the Krylov complexity and the Krylov entropy. The Krylov complexity is determined by the coefficients $r_{k}$ and the phases $\phi_{k}$, which are in turn governed by the Hamiltonian operator given in Eq. (35).
V.1 Krylov complexity
@@
To obtain the Krylov complexity, the key quantity is the wave function. Specifically, the two-mode squeezed state wave function is expressed in the Krylov basis as follows,
$\mathcal{O}(\eta)=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(\eta)|\mathcal{O}_{n})=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}},$ (40)
where $\mathcal{O}(\eta)$ is the operator of wave function whose original definition can be found in Eq. (8). Then, we find the explicit wave function $\phi_{n}$
@@
We define the Krylov compplexity in terms of $\phi_{n}$
@@
This result explicitly demonstrates the independence of the Krylov complexity from $\phi_{k}$; hence, only $r_{k}$ is plotted.
@@
Based on Eq. (42), we show its numerical evolution in Fig. 4, which depicts the behavior of the Krylov complexity throughout the inflationary, RD, and MD eras. Fig. 4 clearly reveals a significant growth in the Krylov complexity during the inflationary phase, a result consistent with the modified dispersion case studied in Li:2024kfm . This finding is also in agreement with related work exploring Krylov complexity under exponential (de Sitter) expansion corresponding to inflation Adhikari:2022oxr . We note that Ref. Adhikari:2022oxr identifies the Krylov complexity with the average particle number, which scales proportionally with volume. However, we emphasize that the approach adopted in Adhikari:2022oxr is specifically applicable to a closed system.
During the RD and MD periods, the Krylov complexity is observed to increase toward certain constant values, around which it subsequently fluctuates. Furthermore, different inflationary models exhibit broadly similar trends in the evolution of Krylov complexity across the respective epochs. Our results are consistent with the analysis in Barbon:2019wsy , which found that in a closed system the Krylov complexity saturates to approximate constant values even after the scrambling time. The scrambling time itself will be analyzed in the open-system framework, specifically in relation to the Lanczos coefficients, the Lyapunov exponent, among other quantities.
V.2 Krylov Entropy
In quantum systems, entropy is defined as a measure of a system’s disorder. The Krylov entropy (K-entropy) serves to quantify the degree of disorder in curvature perturbations. Following Barbon:2019wsy , the K-entropy is defined as follows,